# DyStream: Streaming Dyadic Talking Heads Generation via Flow Matching-based Autoregressive Model

### v1

@@ 5 Limitations and Future Works
  • Metric Definitions (Section D).
@@ Appendix B Disentangled AE for Motion Representation
The VAE is an image-level autoencoder that disentangles a source image $I_{s}$ into an appearance feature $\mathbf{v}_{app}$ and a motion feature $\mathbf{m}$ in the latent space. To enhance the robustness of facial representations, we leverage a 3D-aided appearance encoder ($E_{app}$) and a face decoder ($D_{vae}$) with the same network architecture as LivePortrait [guo2024liveportrait]. The 3D appearance feature volume provides a more accurate characterization of facial appearance in 3D compared to traditional 2D feature maps. We also employ a 2D-based motion encoder ($E_{m}$) to extract motion codes and compress them into low-dimensional 1D representations. This design effectively captures the core semantics of facial motion while avoiding entanglement with appearance information [tishby2000TheInformation, wang2023progressive, wang2024lia].
@@
Readers are referred to [wang2024lia] for more details of this architecture.
@@
The VAE is trained with a composite loss function. We adopt the base losses from LIA [wang2024lia], which include reconstruction ($\mathcal{L}_{rec}$), perceptual ($\mathcal{L}_{feat}$), and adversarial ($\mathcal{L}_{adv}$) losses. We adopt two additional losses following [zakharov2022megaportraits]: a cycle consistency loss ($\mathcal{L}_{cyc}$) to prevent appearance leakage and enhance the disentanglement between motion and appearance features; and a gaze direction loss ($\mathcal{L}_{gaze}$) to encourage diversity in eye movement. To further enhance the disentanglement between motion and appearance features, we follow MegaPortrait [zakharov2022megaportraits] and incorporate severl additional losses: a cross-identity consistency loss ($\mathcal{L}_{cyc}$), an identity consistency loss ($\mathcal{L}_{id}$) a gaze direction loss ($\mathcal{L}_{gaze}$). We additionally incorporate a gaze direction loss ($\mathcal{L}_{gaze}$) to encourage diversity in eye movement. The final objective is a weighted sum of these terms:
@@ Appendix C Inference Performance
Furthermore, we conduct tests on a consumer-grade host equipped with a consumer-grade GPU (4090) and an i9-13900KF CPU. On this setup, our method achieves a FPS of 58.48 and an APD of 0.0171 with a flow matching step of 5. The entire process consumed only 2.37G of VRAM. This is mainly because our inference pipeline is latency-bound and involves frequent CPU–GPU synchronization, where high CPU single-thread performance plays a crucial role. As our current implementation already meets real-time performance requirements, we focus the main paper on the modeling and algorithmic rather than extensive system-level optimization. We note that additional engineering efforts, such as reducing CPU–GPU offloading and synchronization during inference, could further improve runtime performance, and we leave these optimizations to future work.
Appendix D Metric Definitions
@@
We employ the confidence score derived from SyncNet [Chung16a_SyncNet], referred to as SyncScore, to evaluate the temporal synchronization quality between the generated lip movements and the corresponding audio signals. Specifically, we extract audio features and visual features from consecutive video frames, and compute the cross-correlation between audio and visual embeddings across multiple temporal shifts. The confidence score represents the maximum correlation value, indicating the degree of audio-visual synchronization. The final metric is computed as a weighted average across all frames, where higher values indicate better lip-sync quality.
@@
To quantify the presence of motion in generated videos, we utilize the Dynamic Degree metric based on optical flow analysis. We employ the RAFT model [raft] to estimate inter-frame optical flow, sampling frames at intervals of fps/8. For each consecutive frame pair, we compute the flow magnitude and take the mean of the top 5% largest magnitudes as the motion score. A video is classified as dynamic if the number of frames exceeding an adaptive threshold (scaled by resolution and video length) surpasses a minimum count threshold. The final metric represents the proportion of videos exhibiting significant motion.
@@
Following L2L [ng2022learning2listen], this metric quantifies the magnitude and dynamism of generated facial motions by calculating the variance along the temporal dimension. We extract facial motion representations using MediaPipe Face Landmarker, obtaining 6-dimensional head motion parameters (3D translation and 3D rotation in Rodrigues form) and 52-dimensional expression blendshapes for each frame. The variance is computed separately for head motion ($\text{Var}_{\text{head}}$) and facial expressions ($\text{Var}_{\text{exp}}$) as $\text{Var}=\sum_{d=1}^{D}\text{Var}(\mathbf{x}^{(d)})$, where $\mathbf{x}^{(d)}$ denotes the motion sequence along dimension $d$. Higher variance indicates more dynamic and expressive motions.
@@
Following L2L [ng2022learning2listen], this metric quantifies the magnitude and dynamism of generated facial motions by calculating the variance along the temporal dimension. We extract facial motion representations using MediaPipe Face Landmarker, obtaining 6-dimensional head motion parameters (3D translation and 3D rotation in Rodrigues form) and 52-dimensional expression blendshapes for each frame. The variance is computed separately for head motion ($\text{Var}_{\text{head}}$) and facial expressions ($\text{Var}_{\text{exp}}$) as $\text{Var}=\sum_{d=1}^{D}\text{Var}(\mathbf{x}^{(d)})$, where $\mathbf{x}^{(d)}$ denotes the motion sequence along dimension $d$. Higher variance indicates more dynamic and expressive motions.

### v2

@@ 5 Limitations and Future Works
  • Metric Definitions (Section E).
@@ Appendix B Disentangled AE for Motion Representation
The VAE is an image-level autoencoder that disentangles a source image $I_{s}$ into an appearance feature $\mathbf{v}_{app}$ and a motion feature $\mathbf{m}$ in the latent space. To enhance the robustness of facial representations, we leverage a 3D-aided appearance encoder ($E_{app}$) and a face decoder ($D_{vae}$) with the same network architecture as LivePortrait [18]. The 3D appearance feature volume provides a more accurate characterization of facial appearance in 3D compared to traditional 2D feature maps. We also employ a 2D-based motion encoder ($E_{m}$) to extract motion codes and compress them into low-dimensional 1D representations. This design effectively captures the core semantics of facial motion while avoiding entanglement with appearance information [59, 62].
@@
Readers are referred to [62] for more details of this architecture.
@@
The VAE is trained with a composite loss function. We adopt the base losses from LIA [62], which include reconstruction ($\mathcal{L}_{rec}$), perceptual ($\mathcal{L}_{feat}$), and adversarial ($\mathcal{L}_{adv}$) losses. We adopt two additional losses following [13]: a cycle consistency loss ($\mathcal{L}_{cyc}$) to prevent appearance leakage and enhance the disentanglement between motion and appearance features; and a gaze direction loss ($\mathcal{L}_{gaze}$) to encourage diversity in eye movement. To further enhance the disentanglement between motion and appearance features, we follow MegaPortrait [13] and incorporate severl additional losses: a cross-identity consistency loss ($\mathcal{L}_{cyc}$), an identity consistency loss ($\mathcal{L}_{id}$) a gaze direction loss ($\mathcal{L}_{gaze}$). We additionally incorporate a gaze direction loss ($\mathcal{L}_{gaze}$) to encourage diversity in eye movement. The final objective is a weighted sum of these terms:
@@ Appendix C Inference Performance
Furthermore, we conduct tests on a consumer-grade host equipped with a consumer-grade GPU (4090) and an i9-13900KF CPU. On this setup, our method achieves a FPS of 58.48 and an APD of 0.0171. The entire process consumed only 2.37G of VRAM.
Appendix D Head Pose Stability (Anchor)
To quantitatively evaluate the anti-drifting efficacy, we measure the 6-DoF head pose drift relative to the initial frame using MediaPipe. We report Mean/Max Drift (average/maximum deviation) and Std Drift (fluctuation stability). As shown in Table 5, we tested on 10 audio clips (3 minutes each) driven by 10 different reference images. Our method demonstrates exceptional stability across all durations (10s, 60s, 180s). Specifically, our rotation drift remains under $8^{\circ}$ even after 3 minutes, significantly outperforming the “Random Anchor” and “No Anchor” baselines, which suffer from severe cumulative error (drifting $34^{\circ}$).
Table 5: Head Pose Drift Evaluation. We evaluate head pose stability across three different anchor strategies. The metrics reported for Rotation are in degrees, and for Position are in relative units from MediaPipe. Lower values indicate better stability.
Duration Method Rotation (degrees) Position
Mean Drift $\downarrow$ Max Drift $\downarrow$ Std Drift $\downarrow$ Mean Drift $\downarrow$ Max Drift $\downarrow$ Std Drift $\downarrow$
10s Ours $\mathbf{6.61}$ $\mathbf{12.48}$ $\mathbf{2.25}$ $\mathbf{1.04}$ $\mathbf{2.45}$ $\mathbf{0.53}$
Random Anchor $11.65$ $20.76$ $3.83$ $1.62$ $3.32$ $0.74$
No Anchor $30.12$ $41.08$ $8.47$ $2.19$ $4.17$ $0.84$
60s Ours $\mathbf{7.73}$ $\mathbf{17.61}$ $\mathbf{2.45}$ $\mathbf{1.16}$ $\mathbf{3.26}$ $\mathbf{0.57}$
Random Anchor $12.69$ $24.27$ $3.49$ $1.62$ $3.87$ $0.76$
No Anchor $33.96$ $43.25$ $4.59$ $2.38$ $4.56$ $0.73$
180s Ours $\mathbf{7.62}$ $\mathbf{19.31}$ $\mathbf{2.72}$ $\mathbf{1.20}$ $\mathbf{3.69}$ $\mathbf{0.59}$
Random Anchor $12.86$ $26.90$ $4.62$ $1.62$ $4.09$ $0.75$
No Anchor $34.52$ $44.45$ $3.31$ $2.42$ $4.68$ $0.71$
Appendix E Metric Definitions
@@
We employ the confidence score derived from SyncNet [8], referred to as SyncScore, to evaluate the temporal synchronization quality between the generated lip movements and the corresponding audio signals. Specifically, we extract audio features and visual features from consecutive video frames, and compute the cross-correlation between audio and visual embeddings across multiple temporal shifts. The confidence score represents the maximum correlation value, indicating the degree of audio-visual synchronization. The final metric is computed as a weighted average across all frames, where higher values indicate better lip-sync quality.
@@
To quantify the presence of motion in generated videos, we utilize the Dynamic Degree metric based on optical flow analysis. We employ the RAFT model [54] to estimate inter-frame optical flow, sampling frames at intervals of fps/8. For each consecutive frame pair, we compute the flow magnitude and take the mean of the top 5% largest magnitudes as the motion score. A video is classified as dynamic if the number of frames exceeding an adaptive threshold (scaled by resolution and video length) surpasses a minimum count threshold. The final metric represents the proportion of videos exhibiting significant motion.
@@
Following L2L [42], this metric quantifies the magnitude and dynamism of generated facial motions by calculating the variance along the temporal dimension. We extract facial motion representations using MediaPipe Face Landmarker, obtaining 6-dimensional head motion parameters (3D translation and 3D rotation in Rodrigues form) and 52-dimensional expression blendshapes for each frame. The variance is computed separately for head motion ($\text{Var}_{\text{head}}$) and facial expressions ($\text{Var}_{\text{exp}}$) as $\text{Var}=\sum_{d=1}^{D}\text{Var}(\mathbf{x}^{(d)})$, where $\mathbf{x}^{(d)}$ denotes the motion sequence along dimension $d$. Higher variance indicates more dynamic and expressive motions.
@@
Following L2L [42], this metric quantifies the magnitude and dynamism of generated facial motions by calculating the variance along the temporal dimension. We extract facial motion representations using MediaPipe Face Landmarker, obtaining 6-dimensional head motion parameters (3D translation and 3D rotation in Rodrigues form) and 52-dimensional expression blendshapes for each frame. The variance is computed separately for head motion ($\text{Var}_{\text{head}}$) and facial expressions ($\text{Var}_{\text{exp}}$) as $\text{Var}=\sum_{d=1}^{D}\text{Var}(\mathbf{x}^{(d)})$, where $\mathbf{x}^{(d)}$ denotes the motion sequence along dimension $d$. Higher variance indicates more dynamic and expressive motions.